OpenStudy (anonymous):

multiply and simplify √3x(√27x-2√3) the 27x is under the first √

OpenStudy (anonymous):

√a √b = √ab

OpenStudy (jiteshmeghwal9):

\[\sqrt{3x}(\sqrt{27x})-\sqrt{3x}(2\sqrt{3})=?\]

OpenStudy (anonymous):

√81x2-2√9x

OpenStudy (jiteshmeghwal9):

right :)

OpenStudy (anonymous):

9x-2*3x 9x-6x 3x

OpenStudy (jiteshmeghwal9):

right:)

OpenStudy (anonymous):

ok thank you......